# PrivTuner with Homomorphic Encryption and LoRA: A P3EFT Scheme for Privacy-Preserving Parameter-Efficient Fine-Tuning of AI Foundation Models

### v1

@@ V-A Consumption on Mobile Devices
Encryption Time. Let $D_{n}$ be the total number of samples in data $X_{n}$, and $s_{n}$ be the number of parameters per sample that need to be encrypted. Then the encryption time $T_{n}^{\textnormal{en}}$ is:
@@
where $y_{1}(\lambda_{n})$ denotes the required CPU cycles for encrypting per parameter under polynomial modulus $\lambda_{n}$, and $g_{n}$ is the computing capacity on device $n$. The specific expression of $y_{1}(\lambda_{n})$ will be discussed in Sec. VI-B.
@@
For the transmission periods, since the ciphertext data $\widetilde{X}_{n}^{\textnormal{ft}}$ is usually much larger than the plaintext loss $L_{n}$, we focus on the transmission of the previous one. Following [9], we adopt Frequency Division Multiple Access (FDMA) for wireless transmission. In this context, the total available bandwidth is represented as $B^{\textnormal{total}}$. The bandwidth allocated to user $n$ is $B_{n}$ and we have $\sum_{n=1}^{N}B_{n}\leq B^{\textnormal{total}}$. According to the Shannon formula, we obtain the transmission rate $r_{n}$ from user $n$ to the server as follows:
@@
$E_{n}^{\textnormal{tr}}=p_{n}T_{n}~{}\forall n\in\mathcal{N}.$ (14)
V-B Consumption on the Server
The server in a network often houses abundant communication resources. Hence, our focus is primarily on computational consumption, i.e., the prediction process in Step 3 and the adapter updates in Step 5 of Algorithm 1.
@@ V-B Consumption on the Server
The major operations involved in model prediction are addition and multiplication, accompanied by certain other operations, e.g., rotation and bootstrapping. Hence, the total number of CPU cycles $C^{\textnormal{total}}$ is measured by:
$C^{\textnormal{total}}=C^{\textnormal{add}}+C^{\textnormal{mul}}+C^{\textnormal{other}},$ (15)
where $C^{\textnormal{add}}$ accounts for the CPU cycles of all addition operations, $C^{\textnormal{mul}}$ for all multiplication operations, and $C^{\textnormal{other}}$ for cycles required by other operations.
Furthermore, given that the server is aware of the model structure and input dimensionality [^2] For instance, inputs to the $\mathrm{BERT}_{\textnormal{base}}$ model are usually padded or truncated to $X^{\textnormal{input}}\in\mathbb{R}^{l\times d}$, where $l$ denotes the default token length (128, 256, or 512) and $d$ represents the dimensionality (768)., it can calculate the total number of element-wise additions and multiplications required for prediction. In FHE, additions and multiplications are applied component-wisely [20]. Hence, with $a_{n}$ and $m_{n}$ representing the count of additions and multiplications needed per sample during prediction, we estimate CPU cycles as follows:
$y_{2}(\lambda_{n})=y_{3}(\lambda_{n})a_{n}+y_{4}(\lambda_{n})m_{n}+C_{n}^{\textnormal{other}},~{}\forall n\in\mathcal{N},$ (16)
where $y_{2}(\lambda_{n})$ denotes the total CPU cycles to predict one sample under polynomial degree $\lambda_{n}$. $y_{3}(\lambda_{n})$ and $y_{4}(\lambda_{n})$ correspond to cycles for each addition and multiplication. $C_{n}^{\textnormal{other}}$ represents CPU cycles for other operations. The expressions of $y_{3}(\lambda_{n})$ and $y_{4}(\lambda_{n})$ are discussed in Sec. VI-B.
@@
$T_{n}^{\textnormal{cmp}}=\frac{\big{(}y_{2}(\lambda_{n})+c_{n}\big{)}D_{n}}{f_{n}},~{}\forall n\in\mathcal{N},$ (17)
@@
$E_{n}^{\textnormal{cmp}}=\kappa\cdot\big{(}y_{2}(\lambda_{n})+c_{n}\big{)}D_{n}f_{n}^{2},~{}\forall n\in\mathcal{N}.$ (18)
@@ V-C Privacy Protection Level
$\mathcal{S}_{n}=y_{5}(\lambda_{n}),~{}\forall n\in\mathcal{N},$ (19)
where $y_{5}(\lambda_{n})$ is discussed in Sec. VI-B. In our study, we also consider that different devices could have different privacy concerns. With $\sigma_{n}$ serving as a weight parameter denoting the concern for the privacy protection of device $n$, the overall privacy level is quantified as the sum of the individual ones of each device:
@@ VI-A The Studied Optimization Problem
$\mathbb{P}_{0}:\min_{\bm{f},\bm{g},\bm{p},\bm{B},\bm{\lambda}}\left\{\sum_{n=1}^{N}(E_{n}^{\textnormal{en}}\!+\!E_{n}^{\textnormal{tr}}\!+\!E_{n}^{\textnormal{cmp}})\!-\!\omega\sum_{n=1}^{N}\sigma_{n}\mathcal{S}_{n}\right\},$ (23)
@@
where $\bm{f},\bm{g},\bm{p},\bm{B},\bm{\lambda}$ are optimization variables, and $\omega$ in (23) is the weight parameter to balance the energy consumption and privacy protection. Constraints (23a) and (23c) limit the available computation frequencies allocated to device $n$ at the server and device $n$, respectively. Constraints (23b) and (23d) limit the available communication resources, i.e., transmission power and bandwidth. Constraint (23e) provides $m$ choices for the polynomial modulus degrees: $\{\lambda_{\textnormal{o}1},\ldots,\lambda_{\textnormal{o}m}\}$, and ranks them in ascending order. Constraints (28a) and (28b) set the maximum time budgets for both devices and the server.
@@ VI-B Expressions of Functions Used in the Optimization Problem
To conduct a comprehensive analysis of the solution to $\mathbb{P}_{0}$, it is essential to specify the expressions of functions $y_{1}(\lambda_{n})$, $y_{3}(\lambda_{n})$, $y_{4}(\lambda_{n})$, and $y_{5}(\lambda_{n})$ (note that $y_{2}(\lambda_{n})$ uses $y_{3}(\lambda_{n})$ and $y_{4}(\lambda_{n})$, as given in (16)). We first establish the following Condition 1 for functions $y_{1}(\lambda_{n})$, $y_{3}(\lambda_{n})$ and $y_{4}(\lambda_{n})$:
@@
The functions $y_{1}(\lambda_{n})$, $y_{3}(\lambda_{n})$ and $y_{4}(\lambda_{n})$ are all convex with respect to the variable $\lambda_{n}$.
@@
This assertion of convexity is based on empirical observations in [55] as well as our experiments in Section VIII. Convex forms simplify analysis and show satisfactory performance.
@@
$y_{1}(\lambda_{n})=C_{1}(\lambda_{n}+C_{2})^{2},~{}\forall n\in\mathcal{N},$ (24)
where $C_{1},C_{2}>0$ are constants obtained from the curve fitting experiment, to be elaborated in Section VIII.
@@
$y_{3}(\lambda_{n})$ $=C_{3}\lambda_{n}+C_{4},~{}\forall n\in\mathcal{N},$ (25)
$y_{4}(\lambda_{n})$ $=C_{5}\lambda_{n}+C_{6},~{}\forall n\in\mathcal{N},$ (26)
where $C_{3},C_{5}>0$, and $C_{4},C_{6}$ are constants derived from curve fitting. Their specific values are also presented in Section VIII.
For function $y_{5}(\lambda_{n})$, we also give the following condition:
@@
The function $y_{5}(\lambda_{n})$ is concave regarding $\lambda_{n}$.
@@
Via experiments in Section VIII, we find constants $C_{7}>0$ and $C_{8}$ to curve-fit the privacy protection function $y_{5}(\lambda_{n})$:
$y_{5}(\lambda_{n})=C_{7}\lambda_{n}+C_{8},~{}\forall n\in\mathcal{N}.$ (27)
@@ VI-C2 Non-Convexity
To address the above challenges, we will introduce an alternating optimization algorithm in the following Section VII.
@@ VII Algorithm Design
Stage 1 in Sec. VII-A optimizes computation energy and privacy protection while transmission-related variables are fixed. Stage 2 in Sec. VII-B further optimizes the remaining variables based on the outcomes from Stage 1. The overall RA algorithm and its analysis are detailed in Sec. VII-C.
@@ VII-A Stage 1: Optimization in Computation Energy and Privacy Protection Level
In Stage 1, we focus on the optimization of $(\bm{f},\bm{g},\bm{\lambda})$. Given fixed $(\bm{p},\bm{B})$, problem $\mathbb{P}_{0}$ could be simplified into $\mathbb{P}_{1}$:
$\mathbb{P}_{1}:\min_{\bm{f},\bm{g},\bm{\lambda}}~{}$ $\sum_{n=1}^{N}\Big{(}\kappa y_{1}(\lambda_{n})D_{n}s_{n}g_{n}^{2}+\kappa\big{(}c_{n}+y_{2}(\lambda_{n})\big{)}D_{n}f_{n}^{2}\Big{)}$
$-\omega\sum_{n=1}^{N}\sigma_{n}y_{5}(\lambda_{n}),$ (28)
@@
$\frac{\big{(}c_{n}+y_{2}(\lambda_{n})\big{)}D_{n}}{f_{n}}\leq T^{\textnormal{max}}_{S},~{}\forall n\in\mathcal{N}.$ (28b)
@@
$\overline{f}_{n}(\lambda_{n})=\frac{\big{(}c_{n}+y_{2}(\lambda_{n})\big{)}D_{n}}{T^{\textnormal{max}}_{S}},~{}\forall n\in\mathcal{N}.$ (29b)
@@
$\mathbb{P}_{2}:\min_{\bm{\lambda}}$ $\sum_{n=1}^{N}\kappa y_{1}(\lambda_{n})D_{n}s_{n}\overline{g}_{n}^{2}(\lambda_{n})-\omega\sum_{n=1}^{N}\sigma_{n}y_{5}(\lambda_{n})$
$+\sum_{n=1}^{N}\kappa\big{(}c_{n}\!+\!y_{2}(\lambda_{n})\big{)}D_{n}\overline{f}_{n}^{2}(\lambda_{n}),$ (31)
@@
3Set =$\mathcal{L}=\{\mathbb{P}_{3}\}$ and $f_{\mathbb{P}}(\lambda^{*})=\infty$;
@@ 6      Call (47) to solve $P(0)\rightarrow\bm{\lambda}^{\prime}$;
7      if $\bm{\lambda}^{\prime}$ satisfies the integrality (23e) and $f_{\mathbb{P}}(\bm{\lambda}^{\prime})<f_{\mathbb{P}}(\bm{\lambda}^{*})$ then
@@ 20                   Fathom $P$ from $\mathcal{L}$;
$g^{*}_{n}=\frac{\big{(}c_{n}+y_{2}(\lambda_{n}^{*})\big{)}D_{n}}{T^{\textnormal{max}}_{S}},~{}\forall n\in\mathcal{N}.$ (33)
@@ VII-B Stage 2: Optimization in Transmission Energy
$\mathbb{P}_{4}:\min_{\bm{p},\bm{B}}~{}$ $\sum_{n=1}^{N}\frac{p_{n}d_{n}}{r_{n}},$ (34)
@@
For Problem $\mathbb{P}_{3}$, we have:
@@
$\mathbb{P}_{5}:\min_{\bm{p}\bm{B}}~{}$ $\sum_{n=1}^{N}\big{(}(p_{n}d_{n})^{2}z_{n}+\frac{1}{4(r_{n})^{2}z_{n}}\big{)},$ (35)
@@
where $\bm{z}:=[z_{1},\ldots,z_{N}]$ is the introduced auxiliary variable. It is worth noting that problem $\mathbb{P}_{5}$ is a convex problem now and can be solved by analyzing its KKT Conditions. The process of utilizing this method to update $\bm{z}$ and solve $\mathbb{P}_{5}$ is listed in Algorithm 3. It starts by calculating the initial values of auxiliary variables $\bm{z}^{(0)}$, then alternatively refines $(\bm{p},\bm{B})$ and $\bm{z}$ until convergence or reaching maximum iteration number $I$.

### v3

@@ V-A Consumption on Mobile Devices
Encryption Time. Let $D_{n}$ be the total number of samples in data $X_{n}^{\textnormal{ft}}$, and $s_{n}$ be the number of tokens per sample that need to be encrypted. Then the encryption time $T_{n}^{\textnormal{en}}$ is:
@@
where $y_{1}(\lambda_{n})$ denotes the required CPU cycles for encrypting per token under the polynomial modulus $\lambda_{n}$, and $g_{n}$ is the allocated computing capacity on device $n$.
@@
For the transmission periods, since the ciphertext data $\widetilde{X}_{n}^{\textnormal{ft}}$ is usually much larger than the plaintext loss $L_{n}$, we focus on the transmission of the previous one. Following [10], we adopt Frequency Division Multiple Access (FDMA) for wireless transmission. In this context, the total available bandwidth is represented as $B^{\textnormal{total}}$. The bandwidth allocated to user $n$ is $B_{n}$ and we have $\sum_{n=1}^{N}B_{n}\leq B^{\textnormal{total}}$. According to the Shannon formula, we could obtain the uplink transmission rate $r_{n}$ from user $n$ to the server as follows:
@@
$E_{n}^{\textnormal{tr}}=p_{n}T_{n}^{\textnormal{tr}},~\forall n\in\mathcal{N}.$ (14)
V-B Consumption on Server
The server in a network often houses abundant communication resources. Hence, our focus is primarily on the computational consumption of the server, i.e., the encrypted prediction process in Step 3 of Algorithm 1.
@@ V-B Consumption on Server
The major operations involved in model prediction are addition, multiplication and rotation, accompanied by other customized operations, e.g., bootstrapping. Hence, the total number of CPU cycles $C^{\textnormal{total}}$ could be measured by:
$C^{\textnormal{total}}=C^{\textnormal{add}}+C^{\textnormal{mul}}+C^{\textnormal{rot}}+C^{\textnormal{other}},$ (15)
where $C^{\textnormal{add}}$ accounts for the CPU cycles of all additions, $C^{\textnormal{mul}}$ for all multiplications, $C^{\textnormal{rot}}$ for all rotations, and $C^{\textnormal{other}}$ for cycles required by other operations.
Since the server knows the model structure and the utilized FHE techniques, it can roughly count the number of additions, multiplications and rotations required for prediction. Hence, with $a_{n}$, $m_{n}$ and $o_{n}$ denoting the count of additions, multiplications, and rotations needed per sample, we estimate CPU cycles as follows:
$y_{2}(\lambda_{n})\!=\!y_{3}(\lambda_{n})a_{n}\!+\!y_{4}(\lambda_{n})m_{n}\!+\!y_{5}(\lambda_{n})o_{n}\!+\!C_{n}^{\textnormal{other}},~\forall n\in\mathcal{N},$ (16)
where $y_{2}(\lambda_{n})$ denotes the total CPU cycles under $\lambda_{n}$. Besides, $y_{3}(\lambda_{n})$, $y_{4}(\lambda_{n})$ and $y_{5}(\lambda_{n})$ correspond to cycles for each addition, multiplication and rotation, respectively. $C_{n}^{\textnormal{other}}$ represents CPU cycles for other operations. The expressions of $y_{3}(\lambda_{n})$, $y_{4}(\lambda_{n})$ and $y_{5}(\lambda_{n})$ are discussed in Sec. VI-B.
@@
$T_{n}^{\textnormal{cmp}}=\frac{y_{2}(\lambda_{n})D_{n}}{f_{n}},~\forall n\in\mathcal{N},$ (17)
@@
$E_{n}^{\textnormal{cmp}}=\kappa y_{2}(\lambda_{n})D_{n}f_{n}^{2},~\forall n\in\mathcal{N}.$ (18)
@@ V-C Privacy Protection Level
$\mathcal{S}_{n}=y_{6}(\lambda_{n}),~\forall n\in\mathcal{N},$ (19)
where $y_{6}(\lambda_{n})$ is discussed in Sec. VI-B. We also consider that different devices could have different privacy concerns. With $\sigma_{n}$ serving as a weight parameter denoting the concern for the privacy protection of device $n$, the overall privacy level is quantified as the sum of the individual ones of each device:
@@ VI-A The Studied Optimization Problem
$\hskip-75.0pt\mathbb{P}_{0}\mathrel{\mathop{\ordinarycolon}}\min_{\bm{f},\bm{g},\bm{p},\bm{B},\bm{\lambda}}\left\{\sum_{n=1}^{N}(E_{n}^{\textnormal{en}}\!+\!E_{n}^{\textnormal{tr}}\!+\!E_{n}^{\textnormal{cmp}})\!-\!\omega\sum_{n=1}^{N}\sigma_{n}\mathcal{S}_{n}\right\},$ (23)
@@
where $\bm{f},\bm{g},\bm{p},\bm{B}$ and $\bm{\lambda}$ are optimization variables. $\bm{f}=[f_{1},\ldots,f_{N}]$ is a vector containing the allocated frequency to each device at the server, $\bm{g}=[g_{1},\ldots,g_{N}]$ contains the frequency on each device, $\bm{p}=[p_{1},\ldots,p_{N}]$ denotes the transmission power of devices, $\bm{B}=[B_{1},\ldots,B_{N}]$ denotes the allocated bandwidth, and $\bm{\lambda}=[\lambda_{1},\ldots,\lambda_{n}]$ represents the employed polynomial degrees. Besides, $\omega$ in (23) is the weight parameter to balance the energy consumption and privacy protection. Constraints (23a) and (23c) limit the available computation frequencies allocated to device $n$ at the server and device $n$, respectively. Constraints (23b) and (23d) limit the available communication resources, i.e., transmission power and bandwidth. Constraint (23e) provides $m$ choices for the polynomial modulus degrees: $\{\lambda_{\textnormal{o}1},\ldots,\lambda_{\textnormal{o}m}\}$, and ranks them in ascending order. Constraints (23f) and (23g) set the maximum time budgets for both devices and the server.
@@ VI-B Expressions of Functions used in the Optimization Problem
To conduct a comprehensive analysis of the solution to $\mathbb{P}_{0}$, it is essential to specify the expressions of functions $y_{1}(\lambda_{n})$, $y_{3}(\lambda_{n})$, $y_{4}(\lambda_{n})$, $y_{5}(\lambda_{n})$, and $y_{6}(\lambda_{n})$ (note that $y_{2}(\lambda_{n})$ uses $y_{3}(\lambda_{n})$, $y_{4}(\lambda_{n})$ and $y_{5}(\lambda_{n})$, as given in (16)). We first establish the following Condition 1 for those functions:
@@
The functions $y_{1}(\lambda_{n})$, $y_{3}(\lambda_{n})$, $y_{4}(\lambda_{n})$, and $y_{5}(\lambda_{n})$ are all convex with respect to the variable $\lambda_{n}$.
@@
This assertion of convexity is based on empirical observations in [66] and our experiments in Section VIII. Convex forms simplify analysis and show satisfactory performance.
@@
$y_{1}(\lambda_{n})=C_{0}(\lambda_{n}+C_{1})^{2},~\forall n\in\mathcal{N},$ (24)
where $C_{0},C_{1}>0$ are constants obtained from the curve fitting experiment, to be elaborated in Section VIII.
@@
$y_{3}(\lambda_{n})$ $=C_{2}\lambda_{n}+C_{3},~\forall n\in\mathcal{N},$ (25)
$y_{4}(\lambda_{n})$ $=C_{4}\lambda_{n}+C_{5},~\forall n\in\mathcal{N},$ (26)
$y_{5}(\lambda_{n})$ $=C_{6}\lambda_{n}+C_{7},~\forall n\in\mathcal{N},$ (27)
where $C_{2},C_{4},C_{5}>0$, and $C_{3},C_{5},C_{7}$ are all constants derived from curve fitting. Their specific values are also presented in Section VIII.
For function $y_{6}(\lambda_{n})$, we also give the following condition:
@@
The function $y_{6}(\lambda_{n})$ is concave regarding $\lambda_{n}$.
@@
Via experiments in Section VIII, we find constants $C_{8}>0$ and $C_{9}$ to curve-fit the privacy protection function $y_{6}(\lambda_{n})$:
$y_{6}(\lambda_{n})=C_{8}\lambda_{n}+C_{9},~\forall n\in\mathcal{N}.$ (28)
@@ VII Algorithm Design
Stage 1 in Section VII-A optimizes computation energy and privacy protection while transmission-related variables are fixed. Stage 2 in Section VII-B further optimizes the remaining variables based on the outcomes from Stage 1. The overall RA algorithm and its analysis are detailed in Section VII-C.
@@ VII-A Stage 1: Optimization in Computation Energy and Privacy Protection Level
In Stage 1, we first focus on the optimization of $(\bm{f},\bm{g},\bm{\lambda})$ to optimize computation energy consumption and privacy protection level. Given fixed $(\bm{p},\bm{B})$, the original problem $\mathbb{P}_{0}$ could be simplified into $\mathbb{P}_{1}$:
$\mathbb{P}_{1}\mathrel{\mathop{\ordinarycolon}}\min_{\bm{f},\bm{g},\bm{\lambda}}$~ $\sum_{n=1}^{N}\Big(\kappa y_{1}(\lambda_{n})D_{n}s_{n}g_{n}^{2}+\kappa y_{2}(\lambda_{n})D_{n}f_{n}^{2}\Big)$
$-\omega\sum_{n=1}^{N}\sigma_{n}y_{6}(\lambda_{n}),$ (29)
@@
$\frac{y_{2}(\lambda_{n})D_{n}}{f_{n}}\leq T^{\textnormal{max}}_{S},~\forall n\in\mathcal{N}.$ (29b)
@@
$\overline{f}_{n}(\lambda_{n})=\frac{y_{2}(\lambda_{n})D_{n}}{T^{\textnormal{max}}_{S}},~\forall n\in\mathcal{N}.$ (30b)
@@
$\mathbb{P}_{2}\mathrel{\mathop{\ordinarycolon}}\min_{\bm{\lambda}}$ $\sum_{n=1}^{N}\kappa y_{1}(\lambda_{n})D_{n}s_{n}\overline{g}_{n}^{2}(\lambda_{n})-\omega\sum_{n=1}^{N}\sigma_{n}y_{5}(\lambda_{n})$
$+\sum_{n=1}^{N}\kappa y_{2}(\lambda_{n})D_{n}\overline{f}_{n}^{2}(\lambda_{n}),$ (32)
@@
3Set $\mathcal{L}=\{\mathbb{P}_{3}\}$ and $f_{\mathbb{P}}(\lambda^{*})=\infty$;
@@ 6 Call (55) to solve $P(0)\rightarrow\bm{\lambda}^{\prime}$;
7 if $\bm{\lambda}^{\prime}$ satisfies (23e) and $f_{\mathbb{P}}(\bm{\lambda}^{\prime})<f_{\mathbb{P}}(\bm{\lambda}^{*})$ then
@@ 20        Fathom $P$ from $\mathcal{L}$;
$g^{*}_{n}=\frac{y_{2}(\lambda_{n}^{*})D_{n}}{T^{\textnormal{max}}_{S}},~\forall n\in\mathcal{N}.$ (34)
@@ VII-B Stage 2: Optimization in Transmission Energy
$\mathbb{P}_{4}\mathrel{\mathop{\ordinarycolon}}\min_{\bm{p},\bm{B}}$~ $\sum_{n=1}^{N}\frac{p_{n}d_{n}}{r_{n}},$ (35)
@@
For problem $\mathbb{P}_{4}$, we have:
@@
$\mathbb{P}_{5}\mathrel{\mathop{\ordinarycolon}}\min_{\bm{p}\bm{B}}$~ $\sum_{n=1}^{N}\big((p_{n}d_{n})^{2}z_{n}+\frac{1}{4(r_{n})^{2}z_{n}}\big),$ (36)
@@
where $\bm{z}\mathrel{\mathop{\ordinarycolon}}=[z_{1},\ldots,z_{N}]$ is the introduced auxiliary variable. It is worth noting that problem $\mathbb{P}_{5}$ is a convex problem now and can be solved by analyzing its KKT Conditions. The process of utilizing this method to update $\bm{z}$ and solve $\mathbb{P}_{5}$ is listed in Algorithm 3. It starts by calculating the initial values of auxiliary variables $\bm{z}^{(0)}$, then alternatively refines $(\bm{p},\bm{B})$ and $\bm{z}$ until convergence or reaching maximum iteration number $I$.